how would the period of oscillation change if the value of gravity were tripled? how much affect would it have...if at all
1
Expert's answer
2019-10-28T11:39:55-0400

The formula for the period of oscillation:


"T=2\pi\sqrt{\frac{l}{g}}"

If the value of gravity were tripled:


"\frac{T'}{T}=\frac{2\pi\sqrt{\frac{l}{3g}}}{2\pi\sqrt{\frac{l}{g}}}=\frac{1}{\sqrt{3}}"